OpenStudy (anonymous):

Can someone please help me with this question? I will fan and medal!! I think I know the answer, but: A sequence has its first term equal to 8, and each term of the sequence is obtained by adding 6 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence?

OpenStudy (anonymous):

f(n)=f(n-1)+6 f(1)=8

OpenStudy (anonymous):

A. f(1) = 6 and f(n) = f(n - 1) + 8; n > 1 B.f(1) = 8 and f(n) = f(n - 1) + 6; n > 1 C.f(1) = 8 and f(n) = f(n - 1) + 6n; n > 1 D. f(1) = 6 and f(n) = f(n - 1) + 8n; n > 1

OpenStudy (anonymous):

B, first term equals 8 and n th term can be simply obtained by adding to previous function in the series

OpenStudy (anonymous):

Oh okay thanks, that's what I thought.

OpenStudy (anonymous):

ie f(1)=8, f(2)=14, f(3)=20 f(3)=f(3-1)+6 f(3)=f(2)+6 f(2)=f(1)+6 f(1)=8

OpenStudy (anonymous):

=)